The differential equation of a family of tangent lines to the parabola y=x^2, given by 2xt=y+t, t being a parameter,4(y-xdy/dx)+(dy/dx)^2=0.
True or false with statement
1
Expert's answer
2020-10-01T14:09:03-0400

The parametric form of the given equation is "x=t, y=t^2."

The equation of any tangent at "t"  is "2xt=y+t^2."

Differentiating, we get "2t=dy\/dx"

Putting this value in the equation of tangent, we have 


"x(dy\/dx)=y+1\/4(dy\/dx)^2"

"4(y-x(dy\/dx))+(dy\/dx)^2=0"

True.